Show your work. (3) Use the method of substitution to solve the following system of equations: 2x + 3y = 3 8x + 9y = 10
Answer #1

Consider the system of equatras 24+39=3-(1) 8nt qy =10 (2) From (11 34 = 3-2u substitute the value of 3y in (2) weget In +3 (